OpenStudy (anonymous):

sec (-pi/12) I'm doing sum and difference formulas and I don't know how to start this one. Will fan and medal

OpenStudy (ineedhelpfast12):

Which one

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you want to first find \(\cos(-\frac{\pi}{12})\) then flip it to get secant

OpenStudy (misty1212):

if you are going to use the "difference formula" you first have to realize that \[\frac{1}{4}-\frac{1}{3}=-\frac{1}{12}\] so \[\frac{\pi}{4}-\frac{\pi}{3}=-\frac{\pi}{12}\]

OpenStudy (misty1212):

then use the formula to compute \[\cos(\frac{\pi}{4}-\frac{\pi}{3})\]

OpenStudy (misty1212):

you know the formula to use ?

OpenStudy (anonymous):

Yes i do

OpenStudy (anonymous):

@misty1212 I didnt get the answer that my book has

OpenStudy (anonymous):

@amistre64 or @freckles sorry to bother you but do you know how to do this?

OpenStudy (amistre64):

what did you end up with, misty explained it the same i would have

OpenStudy (amistre64):

what is our difference formula for cosine?

OpenStudy (anonymous):

I ended up with 1/4 (sqrt2+sqrt6) The book says its sqrt 6-sqrt 2

OpenStudy (amistre64):

sqrts can be a handful to 'rearrange' lets see what we get cos(a-b) = cosa cosb - sina sinb cos(-pi/12) = cos(pi/4) cos(pi/3) - sin(pi/4) sin(pi/3) cos(-pi/12) = sqrt(2)/2 1/2 - sqrt(2)/2 sqrt(3)/2 cos(-pi/12) = sqrt(2)/4 - sqrt(6)/4 sec(-pi/2) = 4/(sqrt(2)-sqrt(6)) this is what im getting so far ....

OpenStudy (anonymous):

I thought there was a plus in the second part of the formula

OpenStudy (amistre64):

.... good catch, there is. so that just makes the bottom a + instead of a -

OpenStudy (amistre64):

i was wondering how the conjugate was gonna turn out ... i feel better now

OpenStudy (amistre64):

4 --------------- sqrt(2) + sqrt(6) 4 sqrt(2)-sqrt(6) --------------- * ------------- sqrt(2) + sqrt(6) sqrt(2)-sqrt(6) 4(sqrt(2)-sqrt(6)) --------------- 2 - 6 sqrt(6) - sqrt(2)

OpenStudy (anonymous):

Why would I put the four on top?

OpenStudy (amistre64):

because sec = 1/cos

OpenStudy (amistre64):

if cos = a/b then sec = b/a

OpenStudy (amistre64):

we found: cos = k/4 so sec = 4/k

OpenStudy (anonymous):

Oh I got it

OpenStudy (amistre64):

the conjugate (this familiar?) takes care of the simplification

OpenStudy (anonymous):

Can you help me with one more?

OpenStudy (amistre64):

not at the moment, need to attend to some real life issues on my end ....

OpenStudy (anonymous):

Sorry, thanks for all the help :D